OpenStudy (anonymous):

When the Gallup Company conducts a telephone survey, they use a random digit dialer to call potential respondents. The probability that they reach a live person is approximately 0.24. a) What is the probability that they do not reach a live person in five phone calls?

OpenStudy (blurbendy):

P(Does not reach a live person) = 1 - P(Live Person) = 1 - 0.24 = 0.76 they didnt not reach a live person 5 times, so (0.76)^5 = 0.2536

OpenStudy (anonymous):

b) What assumption must you make in order to answer part a? this one too?(: maybe?

OpenStudy (blurbendy):

that each call has the same probability